OpenStudy (anonymous):

It takes 1.5hours for Tim to mom the lawn. Linda can mow the same lawn in 2hours. Approximately how will it take John and Linda, work together, to mow the lawn?

OpenStudy (anonymous):

This is a problem relating to harmonic average. Based on frequency (1/time): Tim: 1/1.5 = 2/3 lawn per hour Linda: 1/2 = 1/2 lawn per hour Tim + Linda: 1/2 + 2/3 = 7/6 lawn per hour Time (1/frequency): 6/7 hour per lawn, therefore working together they take 6/7 hour to mow the lawn

OpenStudy (anonymous):

If you have seen the sum of parallel resistors, the formula works for the same reason. 1/Rt = 1/R1 + 1/R2

OpenStudy (anonymous):

what is the final answer?

OpenStudy (anonymous):

6/7 hour

OpenStudy (anonymous):

tell me in minutes if possible please

OpenStudy (anonymous):

thank you adfriedm

OpenStudy (radar):

Depends on how long John takes to mow the lawn. That information was not provided. Only provided what Tim could do the lawn. No solution, or problem is worded wrong.

OpenStudy (radar):

Jovis209 did you copy this problem correctly?

OpenStudy (anonymous):

yeap i did